OpenStudy (anonymous):

sqrt(cos(sin^2(x))) find the derivative?

hartnn (hartnn):

do u know chain rule ?

OpenStudy (anonymous):

kinda I just barely learning it

hartnn (hartnn):

ok, do u know derivatibe of sqrt x ?

OpenStudy (anonymous):

1/xsqrtx

OpenStudy (anonymous):

1/2sqrt(x)

hartnn (hartnn):

yes, so derivative of \(\sqrt{f(x)}=\dfrac{1}{2\sqrt{f(x)}}\times \dfrac{d}{dx}f'(x)\) applied chain rule, got this ?

OpenStudy (anonymous):

ok

hartnn (hartnn):

so, \(\sqrt{\cos(\sin^2x)}=\dfrac{1}{2\sqrt{\cos(\sin^2x)}}\times \dfrac{d}{dx}[\cos(\sin^2x)]\)

hartnn (hartnn):

so now could you find d/dx of cos sin^2 x ?